Abendsegen (Ges–Dur)

Choralsatz: Flurin Caduff / Transposition für Posaunenchor: Norbert Südland (2013)

[figure: 4 4; 4 4; 4 4; mf]

[figure: 3; Fine]

[figure: mf; mf]

[figure]

[figure] [figure] [figure]

3 mf [figure] [figure] 3 [figure] _{f}

Da Capo al Fine